# Buy versus Build an LLM: A Decision Framework for Governments

### v1

@@ 3.1 Categories of Language Models
  • Monolingual vs. Multilingual. For multi-lingual societies, the breadth of language support is a matter of inclusive governance and equitable service delivery. While buying provides immediate access to globally dominant languages, commercial models might often under-perform in minority languages or dialects due to data scarcity. Building allows governments to prioritize their specific linguistic landscape, though it needs the costly curation of low-resource data. The gap of multilingual support has motivated region-specific efforts such as the SEA-LION family of LLMs [63], which has been developed explicitly to support a wide range of Southeast Asian low-resource languages, including Filipino, Burmese, Lao, Khmer, Tamil and others which are poorly covered by many global models. Similarly, Indian initiatives led by AI4Bharat and Sarvam AI have produced Indic-focused models such as Airavata [28], a Hindi instruction-tuned LLM and Sarvam-1 [4], a multilingual model tailored to major Indian languages, further illustrating how locally developed LLMs can better serve diverse linguistic landscapes than generic global models.
@@
  • Text-only vs. Multi-modal. Multi-modal models expand a system’s input space to include image, audio, and video, allowing for more comprehensive data processing. A nation’s culture is usually captured in text, images, sounds and videos. However, incorporating more modalities adds technical overhead by increasing architectural complexity and resource demands.
@@
  • Reasoning vs. Non-reasoning. Reasoning and non-reasoning models differ in their capacity for long thinking. Compared with non-reasoning models, reasoning models utilize more test-time compute to explore various paths and strategies before producing a final answer [70, 32]. This advanced capability significantly improves performance on complex tasks that require deep deliberation, such as legal interpretation or strategic planning. However, these benefits come with higher computation costs. Governments must weigh the necessity of such depth against the speed of standard models. An emerging alternative is the use of adaptive systems that dynamically allocate a thinking budget or toggle reasoning capabilities based on the complexity of the specific request. [71, 90, 109, 52, 107].
@@
  • Conversational vs. Agentic. Conversational models produce text or media responses within a dialogue. Agentic models go further by interacting with external environments to execute tasks. These agents can perform functions such as browsing websites or databases, calling external tools or managing resource allocation [111, 77]. While agentic systems automate complex workflows and boost productivity, they also introduce significant security risks [30, 27]. Granting models the power to act can lead to unintended system changes or data breaches. Governments must therefore decide how much autonomy to allow. Agentic workflows can be deployed with strict human-in-the-loop controls, approvals, and audit mechanisms, balancing efficiency gains against residual security risk.
@@
  • General-purpose vs. Domain-Specific. General purpose models provide broad utility but often lack the precision required for specialized tasks like legal interpretation or regulatory compliance. Domain specific models focus on narrower applications, such as in healthcare or in finance.   This specialization avoids the cost of training on irrelevant data and reduces hallucinations caused by conflicting information [44, 29]. However, success depends on high quality data and domain expertise, which remain significant hurdles for many government organizations.
@@
  • Open-source vs. Closed-source. The degree of openness varies across different levels, including model parameters, training code, and source datasets. For an open-source model, governments can leverage global community contributions for testing, security auditing, and continuous upgrades. This transparency often builds public trust. However, the potential risks include malicious fine-tuning [34] or unauthorized use for illegal purposes [10]. Moreover, making pre-training data open-source undermines the unique advantage of that country’s models.  In contrast, a closed-source model allows the government to maintain total control over its intellectual property and security perimeter. While this closed-source option protects sensitive architectural details, it requires the government to handle all maintenance and improvements internally without the benefit of community innovation.
@@
  • Model Size. Model size is a primary driver of both capital and operational costs. For those considering a build strategy, larger models require significant capital for high performance hardware and specialized talent. Small models are more accessible for internal development and can be optimized on modest infrastructure. When evaluating a buy strategy, cost is typically tied to usage volume. Larger models often command higher pricing per token, which can lead to high variable costs for frequent tasks. Governments must consider whether the extensive capabilities of a massive model justify the recurring fees or the substantial initial investment required for a private build. In contrast, smaller models may provide sufficient performance for specific administrative tasks while keeping both procurement and maintenance costs manageable.
@@ 3.5 Technical and Resource Demands of Building
  • Expertise. The technical success of a sovereign build depends on a rare combination of expertise across multiple domains including data engineering, infrastructure systems, model training and red-teaming. As established in 2025 labor reports, the global supply of these specialized professionals is exceptionally thin, often leading to “talent wars” where top tier engineers command compensation packages that rival professional athletes [48]. For any institution aiming to build LLMs, the challenge is not just the initial hire, but the retention of institutional memory: the departure of a single lead architect can delay a frontier model project by months or even years. Beyond internal technical expertise, a vibrant domestic open-source ecosystem is also an important resource. A mature developer community acts as a force multiplier which provides a network that can rapidly stress test architectures, build derivative tools, and accelerate the localization of frontier technology.
@@
  • Data. The effectiveness of any model is fundamentally limited by the quality and composition of its training data. This process requires a massive effort to collect, clean, and annotate information while ensuring broad domain coverage. Developers must also master a precise mixing recipe that balances diverse data sources to ensure the model gains general reasoning abilities without losing specialized knowledge. As a reference, the Llama 3 technical report [26] describes a massive pre-training dataset of over 15 trillion tokens, where the data mixture was meticulously tuned to balance code, multilingual content, and high-quality reasoning data. Similarly, DeepSeek-V3 [54] details a strategy involving $14.8$ trillion tokens and a post-training phase that utilizes knowledge distillation to refine the "recipe" for specific task capabilities. Furthermore, data requirements shift throughout the training life-cycle; pre-training requires trillions of diverse tokens to build a knowledge base, while mid-training and post-training demand highly curated datasets for instruction following and safety alignment, a strategy emphasized in the Gemma 2 [89] report which prioritized data quality and distillation over sheer volume. We provide detailed statistics on training data and token counts in Table 1, compiled from publicly available model reports.
@@
  • Architectural Design Choices. A sovereign build also requires governments to make a series of early architectural design choices that often require deep, hard-earned experience. Decisions regarding model architecture, context length, tokenizer, and vocabulary size, as well as modality support, training techniques (e.g., full precision or mixed precision) and objective functions, all these design questions are rarely solved analytically and require extensive experience. For a team without prior large model training experience, this architectual “design tax” can be as significant, and should therefore be treated as a first-order consideration in any decision to build. We compile representative architectural specifications and training costs from several open-source LLM reports and summarize them in Table 1.
@@
  • Training Compute Demands. The physical infrastructure required for training is substantial and expensive. For reference, training a mid-scale model with $70$ billion parameters, such as Llama 3 70B, can require approximately $6.4$ million GPU hours [92], while the subsequent Llama-3.3 70B utilized roughly $7.0$ million GPU hours for its pre-training phase [65]. The architectural choice and training technique adoption can make a difference: as reported in DeepSeek-V3 report [54], training a 671671B DeepSeekMoE architecture with the support of FP8 training requires only $2.8$ million H800 GPU hours for its full training. Additionally, managing these runs requires robust training frameworks capable of handling checkpointing and failure recovery. At this scale, even a brief hardware interruption can result in million dollar losses in wasted compute time. Estimating the training cost must account for unexpected failures and interruptions. For example, Meta has reported in [26] that during the 54-day pretraining of Llama 3 405405B model, more than half of the $419$ unexpected interruptions were caused by failures related to GPUs or their onboard HBM3 memory. Additionally, many of the most effective training techniques remain proprietary or poorly documented, which increases the risks for teams without extensive experience.
@@
  • Evaluation. Evaluation measures the operational utility of a model across various dimensions, including instruction following ability [42, 104, 102], domain-specific proficiency (e.g. coding [41] and math [53, 9]), context window reliability [21], efficiency and so on. Beyond relying on generic leaderboards, governments developing sovereign models must design evaluation frameworks that reflect national priorities rather than purely academic or commercial benchmarks. In practice, governments often have to decide whether to build their own benchmarks: an example was the case in SEA-LION [63] where there were no high-quality evalution datasets for Lao, Khmer, or regional multi-turn dialogues in Malay and Tagalog, requiring the creation of new benchmarks such as SEA-HELM [84]. Constructing sovereign benchmarks brings its own challenges. Localizing test content requires extensive human expertise and cannot be easily automated or synthetically generated at scale - and the use of LLM-as-a-judge methods introduces a “chicken-and-egg” problem.
@@
  • Alignment, Regulation and Red-Teaming. Once a model is trained, it must undergo rigorous alignment to ensure its behavior is safe, lawful and consistent with national or regional norms. This involves constraining the model to reduce the likelihood of generating harmful, misleading, or non-compliant outputs, often requiring constructing specific datasets and evaluation sets, which are time-consuming and manpower-intensive. In addition, robust red-teaming is essential: agencies should proactively stress-test LLMs for adversarial prompting [55, 56], jailbreaking [117, 112] and any other policy-violating behaviors [8]. Systematic red-teaming has become a standard pre-deployment practice in the industry: leading model providers such as OpenAI [3, 15], Anthropic [86] and Google [78] routinely conduct extensive adversarial testing to ensure safety before releasing large models. For government agencies, compliance with sector specific regulations and auditing standards is non-negotiable and a continuous commitment.
@@
  • Inference and Deployment. The final challenge lies in the production environment where scalability and latency are critical. LLMs are expected to process high volumes of concurrent requests while maintaining consistent response speeds. Maintaining operational viability involves balancing computational throughput with cost efficiency to prevent infrastructure overhead from exceeding sustainable levels. Deployment choices also shape governance: production systems need robust safeguards against prompt injection, data exfiltration, and misuse, as well as auditable logging and access controls that meet public-sector requirements. In practice, governments may rely on mature hosting and MLOps platforms to reduce operational burden. For example, managed model hubs and inference services (e.g., Hugging Face), hyperscaler cloud platforms and model marketplaces (e.g., AWS, Azure, Google Cloud), or enterprise AI platforms (e.g., IBM); or adopt hybrid approaches combining on-premise/sovereign-cloud hosting with vendor tools. Ultimately, the practical impact of an AI system depends on a structured roadmap for deployment and tool orchestration that ensures reliability in real-world daily public service delivery.
@@ 4.3 Cost and Financial Perspective
  • Capital Expenditure (CAPEX). A build strategy requires significant upfront investment in physical hardware, such as specialized GPU clusters and high performance networking. It also involves the cost of acquiring massive datasets and hiring a team of elite technical experts. These projects carry higher trial-and-error risks where a failure can lead to substantial financial loss. In contrast, a buy strategy has much lower initial capital requirements, with costs mainly focused on security reviews, network integration and staff training for third-party interfaces.
  • Operating Expenditure (OPEX). For a built model, operating costs are relatively stable but significant, covering energy consumption, hardware maintenance, and permanent staff salaries. However, a volatile and easily overlooked component of OPEX is the cumulative expenditure on experimental overhead - before achieving a final successful training run, the team must fund ablation studies and hyperparameter tuning sessions to identify the optimal model configuration. According to the Frontier Model Training Costs report by Epoch AI [23], the total compute for model development is 1.21.2x to 44x larger than the final training run. For a bought model, costs scale directly with usage volume. While this provides flexibility, it also creates budget uncertainty because vendors can change pricing models or discontinue teaser rates once a department is integrated into their ecosystem. Over time, high volume government use can make a subscription more expensive than the maintenance of an internal system.
@@
  • Revenue and Cost Recovery. Building an internal model allows the government to treat the system as a national asset. The government can recover costs by charging commercial entities or high volume public users for access, effectively reinvesting that revenue into the domestic economy. When buying, fees flow primarily to external vendors.
@@ 4.4 Economic and Regional Development Perspectives
  • National Economic Development Trajectory. The buy–build decision should also be viewed in light of a country’s long-term economic development strategy. Governments pursuing a transition toward a technology- and knowledge-driven economy may place higher strategic value on building domestic LLM capabilities, as such investments can stimulate high-skilled employment, upstream infrastructure development, and downstream innovation across sectors. By contrast, countries whose economic structures are expected to remain primarily agriculture-, resource-, or extraction-based may rationally prioritize access to AI capabilities as users rather than builders, focusing on cost efficiency and rapid deployment over domestic model development. In this sense, the decision to build or buy is not purely technical or financial, but closely tied to national industrial policy and growth objectives.
  • Regional Strategic Positioning. Even if building domestic LLMs is unlikely to place a country among global AI leaders, it can still confer meaningful regional advantages relative to neighboring or peer countries. Indigenous AI capabilities may strengthen a country’s bargaining position in regional cooperation, reduce reliance on and increase bargaining power with external technology providers, and allow it to set standards within shared markets or institutions. In some cases, serving as an early regional adopter or developer of government-grade AI systems can attract talent, partnerships, and investment, reinforcing the country’s strategic relevance within its region.

### v2

@@ Abstract
The short version of this document is published as an ACM TechBrief at https://dl.acm.org/doi/epdf/10.1145/3797946, and this document is published as an ACM Technology Policy Council white paper at https://www.acm.org/binaries/content/assets/public-policy/buildvsbuyai.pdf.
@@ 3.1 Categories of Language Models
  • Monolingual vs. Multilingual: For multi-lingual societies, the breadth of language support is a matter of inclusive governance and equitable service delivery. While buying provides immediate access to globally dominant languages, commercial models might often under-perform in minority languages or dialects due to data scarcity. Building allows governments to prioritize their specific linguistic landscape, though it needs the costly curation of low-resource data. The gap of multilingual support has motivated region-specific efforts such as the SEA-LION family of LLMs [63], which has been developed explicitly to support a wide range of Southeast Asian low-resource languages, including Filipino, Burmese, Lao, Khmer, Tamil and others which are poorly covered by many global models. Similarly, Indian initiatives led by AI4Bharat and Sarvam AI have produced Indic-focused models such as Airavata [28], a Hindi instruction-tuned LLM and Sarvam-1 [4], a multilingual model tailored to major Indian languages, further illustrating how locally developed LLMs can better serve diverse linguistic landscapes than generic global models.
@@
  • Text-only vs. Multi-modal: Multi-modal models expand a system’s input space to include image, audio, and video, allowing for more comprehensive data processing. A nation’s culture is usually captured in text, images, sounds and videos. However, incorporating more modalities adds technical overhead by increasing architectural complexity and resource demands.
@@
  • Reasoning vs. Non-reasoning: Reasoning and non-reasoning models differ in their capacity for long thinking. Compared with non-reasoning models, reasoning models utilize more test-time compute to explore various paths and strategies before producing a final answer [70, 32]. This advanced capability significantly improves performance on complex tasks that require deep deliberation, such as legal interpretation or strategic planning. However, these benefits come with higher computation costs. Governments must weigh the necessity of such depth against the speed of standard models. An emerging alternative is the use of adaptive systems that dynamically allocate a thinking budget or toggle reasoning capabilities based on the complexity of the specific request. [71, 90, 109, 52, 107].
@@
  • Conversational vs. Agentic: Conversational models produce text or media responses within a dialogue. Agentic models go further by interacting with external environments to execute tasks. These agents can perform functions such as browsing websites or databases, calling external tools or managing resource allocation [111, 77]. While agentic systems automate complex workflows and boost productivity, they also introduce significant security risks [30, 27]. Granting models the power to act can lead to unintended system changes or data breaches. Governments must therefore decide how much autonomy to allow. Agentic workflows can be deployed with strict human-in-the-loop controls, approvals, and audit mechanisms, balancing efficiency gains against residual security risk.
@@
  • General-purpose vs. Domain-Specific: General purpose models provide broad utility but often lack the precision required for specialized tasks like legal interpretation or regulatory compliance. Domain specific models focus on narrower applications, such as in healthcare or in finance.   This specialization avoids the cost of training on irrelevant data and reduces hallucinations caused by conflicting information [44, 29]. However, success depends on high quality data and domain expertise, which remain significant hurdles for many government organizations.
@@
  • Open-source vs. Closed-source: The degree of openness varies across different levels, including model parameters, training code, and source datasets. For an open-source model, governments can leverage global community contributions for testing, security auditing, and continuous upgrades. This transparency often builds public trust. However, the potential risks include malicious fine-tuning [34] or unauthorized use for illegal purposes [10]. Moreover, making pre-training data open-source undermines the unique advantage of that country’s models.  In contrast, a closed-source model allows the government to maintain total control over its intellectual property and security perimeter. While this closed-source option protects sensitive architectural details, it requires the government to handle all maintenance and improvements internally without the benefit of community innovation.
@@
  • Model Size: Model size is a primary driver of both capital and operational costs. For those considering a build strategy, larger models require significant capital for high performance hardware and specialized talent. Small models are more accessible for internal development and can be optimized on modest infrastructure. When evaluating a buy strategy, cost is typically tied to usage volume. Larger models often command higher pricing per token, which can lead to high variable costs for frequent tasks. Governments must consider whether the extensive capabilities of a massive model justify the recurring fees or the substantial initial investment required for a private build. In contrast, smaller models may provide sufficient performance for specific administrative tasks while keeping both procurement and maintenance costs manageable.
@@ 3.5 Technical and Resource Demands of Building
  • Expertise: The technical success of a sovereign build depends on a rare combination of expertise across multiple domains including data engineering, infrastructure systems, model training and red-teaming. As established in 2025 labor reports, the global supply of these specialized professionals is exceptionally thin, often leading to “talent wars” where top tier engineers command compensation packages that rival professional athletes [48]. For any institution aiming to build LLMs, the challenge is not just the initial hire, but the retention of institutional memory: the departure of a single lead architect can delay a frontier model project by months or even years. Beyond internal technical expertise, a vibrant domestic open-source ecosystem is also an important resource. A mature developer community acts as a force multiplier which provides a network that can rapidly stress test architectures, build derivative tools, and accelerate the localization of frontier technology.
@@
  • Data: The effectiveness of any model is fundamentally limited by the quality and composition of its training data. This process requires a massive effort to collect, clean, and annotate information while ensuring broad domain coverage. Developers must also master a precise mixing recipe that balances diverse data sources to ensure the model gains general reasoning abilities without losing specialized knowledge. As a reference, the Llama 3 technical report [26] describes a massive pre-training dataset of over 15 trillion tokens, where the data mixture was meticulously tuned to balance code, multilingual content, and high-quality reasoning data. Similarly, DeepSeek-V3 [54] details a strategy involving $14.8$ trillion tokens and a post-training phase that utilizes knowledge distillation to refine the "recipe" for specific task capabilities. Furthermore, data requirements shift throughout the training life-cycle; pre-training requires trillions of diverse tokens to build a knowledge base, while mid-training and post-training demand highly curated datasets for instruction following and safety alignment, a strategy emphasized in the Gemma 2 [89] report which prioritized data quality and distillation over sheer volume. We provide detailed statistics on training data and token counts in Table 1, compiled from publicly available model reports.
@@
  • Architectural Design Choices: A sovereign build also requires governments to make a series of early architectural design choices that often require deep, hard-earned experience. Decisions regarding model architecture, context length, tokenizer, and vocabulary size, as well as modality support, training techniques (e.g., full precision or mixed precision) and objective functions, all these design questions are rarely solved analytically and require extensive experience. For a team without prior large model training experience, this architectual “design tax” can be as significant, and should therefore be treated as a first-order consideration in any decision to build. We compile representative architectural specifications and training costs from several open-source LLM reports and summarize them in Table 1.
@@
  • Training Compute Demands: The physical infrastructure required for training is substantial and expensive. For reference, training a mid-scale model with $70$ billion parameters, such as Llama 3 70B, can require approximately $6.4$ million GPU hours [92], while the subsequent Llama-3.3 70B utilized roughly $7.0$ million GPU hours for its pre-training phase [65]. The architectural choice and training technique adoption can make a difference: as reported in DeepSeek-V3 report [54], training a 671671B DeepSeekMoE architecture with the support of FP8 training requires only $2.8$ million H800 GPU hours for its full training. Additionally, managing these runs requires robust training frameworks capable of handling checkpointing and failure recovery. At this scale, even a brief hardware interruption can result in million dollar losses in wasted compute time. Estimating the training cost must account for unexpected failures and interruptions. For example, Meta has reported in [26] that during the 54-day pretraining of Llama 3 405405B model, more than half of the $419$ unexpected interruptions were caused by failures related to GPUs or their onboard HBM3 memory. Additionally, many of the most effective training techniques remain proprietary or poorly documented, which increases the risks for teams without extensive experience.
@@
  • Evaluation: Evaluation measures the operational utility of a model across various dimensions, including instruction following ability [42, 104, 102], domain-specific proficiency (e.g. coding [41] and math [53, 9]), context window reliability [21], efficiency and so on. Beyond relying on generic leaderboards, governments developing sovereign models must design evaluation frameworks that reflect national priorities rather than purely academic or commercial benchmarks. In practice, governments often have to decide whether to build their own benchmarks: an example was the case in SEA-LION [63] where there were no high-quality evalution datasets for Lao, Khmer, or regional multi-turn dialogues in Malay and Tagalog, requiring the creation of new benchmarks such as SEA-HELM [84]. Constructing sovereign benchmarks brings its own challenges. Localizing test content requires extensive human expertise and cannot be easily automated or synthetically generated at scale - and the use of LLM-as-a-judge methods introduces a “chicken-and-egg” problem.
@@
  • Alignment, Regulation and Red-Teaming: Once a model is trained, it must undergo rigorous alignment to ensure its behavior is safe, lawful and consistent with national or regional norms. This involves constraining the model to reduce the likelihood of generating harmful, misleading, or non-compliant outputs, often requiring constructing specific datasets and evaluation sets, which are time-consuming and manpower-intensive. In addition, robust red-teaming is essential: agencies should proactively stress-test LLMs for adversarial prompting [55, 56], jailbreaking [117, 112] and any other policy-violating behaviors [8]. Systematic red-teaming has become a standard pre-deployment practice in the industry: leading model providers such as OpenAI [3, 15], Anthropic [86] and Google [78] routinely conduct extensive adversarial testing to ensure safety before releasing large models. For government agencies, compliance with sector specific regulations and auditing standards is non-negotiable and a continuous commitment.
@@
  • Inference and Deployment: The final challenge lies in the production environment where scalability and latency are critical. LLMs are expected to process high volumes of concurrent requests while maintaining consistent response speeds. Maintaining operational viability involves balancing computational throughput with cost efficiency to prevent infrastructure overhead from exceeding sustainable levels. Deployment choices also shape governance: production systems need robust safeguards against prompt injection, data exfiltration, and misuse, as well as auditable logging and access controls that meet public-sector requirements. In practice, governments may rely on mature hosting and MLOps platforms to reduce operational burden. For example, managed model hubs and inference services (e.g., Hugging Face), hyperscaler cloud platforms and model marketplaces (e.g., AWS, Azure, Google Cloud), or enterprise AI platforms (e.g., IBM); or adopt hybrid approaches combining on-premise/sovereign-cloud hosting with vendor tools. Ultimately, the practical impact of an AI system depends on a structured roadmap for deployment and tool orchestration that ensures reliability in real-world daily public service delivery.
@@ 4.3 Cost and Financial Perspective
  • Capital Expenditure (CAPEX): A build strategy requires significant upfront investment in physical hardware, such as specialized GPU clusters and high performance networking. It also involves the cost of acquiring massive datasets and hiring a team of elite technical experts. These projects carry higher trial-and-error risks where a failure can lead to substantial financial loss. In contrast, a buy strategy has much lower initial capital requirements, with costs mainly focused on security reviews, network integration and staff training for third-party interfaces.
  • Operating Expenditure (OPEX): For a built model, operating costs are relatively stable but significant, covering energy consumption, hardware maintenance, and permanent staff salaries. However, a volatile and easily overlooked component of OPEX is the cumulative expenditure on experimental overhead - before achieving a final successful training run, the team must fund ablation studies and hyperparameter tuning sessions to identify the optimal model configuration. According to the Frontier Model Training Costs report by Epoch AI [23], the total compute for model development is 1.21.2x to 44x larger than the final training run. For a bought model, costs scale directly with usage volume. While this provides flexibility, it also creates budget uncertainty because vendors can change pricing models or discontinue teaser rates once a department is integrated into their ecosystem. Over time, high volume government use can make a subscription more expensive than the maintenance of an internal system.
@@
  • Revenue and Cost Recovery: Building an internal model allows the government to treat the system as a national asset. The government can recover costs by charging commercial entities or high volume public users for access, effectively reinvesting that revenue into the domestic economy. When buying, fees flow primarily to external vendors.
@@ 4.4 Economic and Regional Development Perspectives
  • National Economic Development Trajectory: The buy–build decision should also be viewed in light of a country’s long-term economic development strategy. Governments pursuing a transition toward a technology- and knowledge-driven economy may place higher strategic value on building domestic LLM capabilities, as such investments can stimulate high-skilled employment, upstream infrastructure development, and downstream innovation across sectors. By contrast, countries whose economic structures are expected to remain primarily agriculture-, resource-, or extraction-based may rationally prioritize access to AI capabilities as users rather than builders, focusing on cost efficiency and rapid deployment over domestic model development. In this sense, the decision to build or buy is not purely technical or financial, but closely tied to national industrial policy and growth objectives.
  • Regional Strategic Positioning: Even if building domestic LLMs is unlikely to place a country among global AI leaders, it can still confer meaningful regional advantages relative to neighboring or peer countries. Indigenous AI capabilities may strengthen a country’s bargaining position in regional cooperation, reduce reliance on and increase bargaining power with external technology providers, and allow it to set standards within shared markets or institutions. In some cases, serving as an early regional adopter or developer of government-grade AI systems can attract talent, partnerships, and investment, reinforcing the country’s strategic relevance within its region.